### 04 Dec 2023 v1

# GNN2R: Weakly-Supervised Rationale-Providing Question Answering over Knowledge Graphs

Abstract
Most current methods for multi-hop question answering (QA) over knowledge graphs (KGs) only provide final conclusive answers without explanations, such as a set of KG entities that is difficult for normal users to review and comprehend. This issue severely limits the application of KG-based QA in real-world scenarios. However, it is non-trivial to solve due to two challenges: First, annotations of reasoning chains of multi-hop questions, which could serve as supervision for explanation generation, are usually lacking. Second, it is difficult to maintain high efficiency when explicit KG triples need to be retrieved to generate explanations. In this paper, we propose a novel Graph Neural Network-based Two-Step Reasoning model (GNN2R) to solve this issue. GNN2R can provide both final answers and reasoning subgraphs as a rationale behind final answers efficiently with only weak supervision that is available through question-final answer pairs. We extensively evaluated GNN2R with detailed analyses in experiments. The results demonstrate that, in terms of effectiveness, efficiency, and quality of generated explanations, GNN2R outperforms existing state-of-the-art methods that are applicable to this task. Our code and pre-trained models are available at https://github.com/ruijie-wang-uzh/GNN2R.
Index Terms:
Knowledge Graphs, Question Answering, Graph Neural Networks, Language Models
I Introduction
In recent years, knowledge graphs (KGs) have been prominently used as knowledge sources supporting multi-hop question-answering (QA) [1]. An example is given in Fig. 1, where the question “who was born in California and directed a movie starring Michael Keaton” can be answered based on the given KG with Tim Burton. This seems to be a successful application of KG-based QA. However, it is difficult to verify and understand the answer for normal users, given the complexity of navigating question-relevant parts within a large KG, not to mention the technical expertise required for such navigation, e.g., SPARQL [2] and understanding of the schema of the KG [3]. We propose to provide users with a reasoning subgraph consisting of triples from the KG that elucidate how the answer satisfies the semantic constraints specified in the question. As highlighted in the figure, the reasoning subgraph of this question consists of two paths: [(Tim Burton, birthplace, California)], which affirms that Tim Burton was born in California, and [(Tim Burton, director-1, Batman 1989),  [^1] We use the superscript -1 to denote the inverse of a relation. (Batman 1989, cast member, Michael Keaton)], which affirms that Tim Burton directed Batman 1989 starring Michael Keaton. Hence, reasoning subgraphs provide users with a rationale for the final answers. Also, in the spirit of transparent AI, intermediate information in reasoning subgraphs, e.g., the alluded movie in this example, could be useful and important.
However, the generation of reasoning subgraphs is non-trivial due to two challenges: First, annotating ground-truth reasoning subgraphs is expensive and, in low-resource scenarios, even infeasible. It is challenging to train a model that generates reasoning subgraphs based on only annotations of final answers. Second, QA is expected to function online and interactively, with low-latency responses to given questions. However, the generation of subgraphs requires explicit reasoning and time-consuming graph-based operations (e.g., path extraction) over KG triples. Especially considering that the search space could be very large with multi-hop questions and large KGs, it is challenging to maintain high efficiency.
Figure 1: An example of KG-based QA, where relevant entities and relations are highlighted in red and blue. The reasoning subgraph is distinguished by a light-orange background, featuring two reasoning paths in orange and purple.
To achieve rationale-providing QA while handling the above challenges, we propose a novel Graph Neural Network-based Two-step Reasoning model, called GNN2R, which answers questions in two steps: First, a graph neural network (GNN) is used to represent KG entities and given questions in joint embedding spaces where questions are close to their answers. This step efficiently prunes the search space of candidate answers and, consequently, reasoning subgraphs. Second, we employ a language model (LM) to extract reasoning subgraphs based on their semantic similarity to the given question. Three algorithms are proposed for the subgraph extraction and the weakly-supervised fine-tuning of LM in this step.
We summarize our contributions in this paper in the following:
  • We propose a novel QA model—GNN2R that can generate both final answers and reasoning subgraphs efficiently with only weak supervision, also achieving state-of-the-art performance in this task.
  • We propose a novel graph neural network (GNN) that can efficiently reason over KGs to identify candidate answers and prune the search space of reasoning subgraphs.
  • We propose a novel approach with simple and effective algorithms to fine-tune an LM using only weak supervision for the extraction of reasoning subgraphs.
  • We conducted extensive experiments to evaluate the effectiveness and efficiency of the proposed model and examine the quality of generated reasoning subgraphs, which resulted in findings useful for future research on this issue.
II Related Work
We introduce three types of current mainstream QA methods and discuss the potential application of large language models.
Query construction-based methods [4, 5, 6, 7, 8, 9, 10] try to translate given questions into logical queries, e.g., SPARQL [2] and symbolic programs [11], that can be executed over KGs to obtain final answers. They have been studied for a long time and have achieved competitive performance. Also, by instantiating the constructed logical queries, these methods can be potentially used to generate reasoning subgraphs. However, recent methods of this type, especially case-based [6], transfer learning-based [7], and supervised learning-based [10], heavily rely on annotations of logical queries, which makes them inapplicable in low-resource scenarios with only weak supervision available, including the task of this paper, where only final answers are annotated for training.
Retrieval-based methods [12, 13, 14, 15, 16, 17] compute contextual KG representations and leverage distance computations or learned neural classifiers to retrieve answers by the parametric ranking or classification of entities. For example, Sun et al. [14] propose to use a graph neural network [13] to compute contextual representations of KG entities and use these embeddings to classify entities into answers and non-answer entities. These methods are highly efficient. However, as these methods do not explicitly reason over triples, they do not offer explanations. In the first step, our model selects candidate answers from computed embedding spaces, which aligns with this category.
Reasoning-based methods [18, 19, 20, 21, 22, 23] are potentially capable of extracting reasoning subgraphs from KGs. A branch [18, 20] of these methods performs explicit explorations in KGs and generates reasoning chains for given questions. For example, Qiu et al. [20] use a reinforcement learning agent that explores paths in the KG starting from topic entities mentioned in given questions. Another branch [21, 22] performs fuzzy reasoning by computing confidence scores or numerical labels for entities and triples in KGs. For example, Shi et al. [21] propose to compute stepwise scores for relations in KGs to determine how likely a triple is part of the reasoning chain of a given question. These methods are the ones most applicable to the generation of reasoning subgraphs. However, the weak supervision and efficiency issues have not been well addressed. Some methods, such as [19], rely on annotations of reasoning chains, and their mechanisms for explicit KG reasoning, such as [20], are prone to inefficiency. Furthermore, most of these methods can only provide probable reasoning chains with likelihoods that are difficult to comprehend.
The application of large language models (LLMs), in KG-based QA has recently attracted wide attention [24, 25, 26]. For example, Tan et al. [25] find that LLMs have the potential to outperform traditional QA methods, which do not rely on LLMs, in a relaxed setup, where a question is deemed correctly answered if the response from LLMs mentions any valid answer. However, this study also highlights a crucial drawback of LLMs—extracting answers from their textual response of varying length is difficult, especially considering that the answers may be expressed as aliases, acronyms, synonyms, or even in a different language. Also, given the implicit nature of the knowledge in LLMs and the common hallucination issue [27], it is challenging to verify the validity and completeness of answers from LLMs. In contrast, traditional QA methods can provide explicit answers in an unambiguous and uniform manner, and our method is able to extract the rationale behind answers from KGs. Furthermore, Sun et al. [26] show that even the current state-of-the-art LLMs, including ChatGPT [^2] https://platform.openai.com/docs/api-reference and LLaMA (33B) [28], still have difficulty answering simple (one-hop) factual questions generated from templates, which is often considered a solved task with traditional QA methods. Therefore, in the era of LLMs, the contributions of this paper, alongside other methods without relying on LLMs, still hold practical significance.
III Problem Definition
We denote a KG as $\mathcal{G}=(E,R,T)$, where $E$, $R$, and $T$ respectively refer to the entity set, relation set, and triple set of $\mathcal{G}$. A triple $(e_{h},r,e_{t})\in T$ consists of a head entity $e_{h}\in E$, a relation $r\in R$, and a tail entity $e_{t}\in E$. A natural language question posed over $\mathcal{G}$ is denoted as $q$. The answer to $q$ is a set of entities in $E$, denoted as $A_{q}\subseteq E$. The reasoning subgraph of $q$, denoted as $\mathcal{G}_{q}$, is a subgraph of $\mathcal{G}$ that consists of triples linking answers in $A_{q}$ with topic entities mentioned in $q$ in the way that satisfies the semantic constraints specified in $q$. In this paper, only a limited set of question-final answer pairs are available for training. The task is to train a QA model that can retrieve both the answers and the reasoning subgraphs for new and unseen questions.
IV Method: The GNN2R QA Model
GNN2R answers a given question in two steps: a graph neural network (GNN)-based coarse reasoning step and a language model (LM)-based explicit reasoning step. In the first step (cf. Section IV-A), a GNN is trained to represent KG entities and the question in a joint embedding space, where the question is close to its answers. We collect candidate answers via nearest neighbor search. In the second step (cf. Section IV-B), we extract candidate reasoning subgraphs that link the candidate answers to topic entities mentioned in the given question. Then, we fine-tune a pre-trained Sentence Transformer [29] to select the final reasoning subgraph that is the most semantically consistent with the given question.
IV-A Step-I: GNN-based Coarse Reasoning
Figure 2: The GNN-based encoding of the given question and KG in Step-I.
The proposed GNN consists of four modules: a general question encoder, a layerwise question encoder, an attentional message propagation module, and a gated embedding update module, which are all highlighted as colored boxes in Fig. 2. There are $K$ layers of KG encoding, where illustrated is the detailed architecture of the k-th layer. The arrows represent module inputs and outputs. Only considering those dark arrows would be sufficient for understanding the network.
The general question encoder takes $\mathbf{Q}=[\mathbf{w}_{1},\mathbf{w}_{2},...,\mathbf{w}_{l}]$ as input, which is the initial encoding of the question $q$. $\mathbf{Q}$ could be any contextual vector representations, e.g., word2vec [30], Glove [31], or BERT encodings [32]. In our implementation, we first use OpenNER [^3] https://github.com/zmd971202/OpenNER [33] to identify mentions of topic entities and mask them out from $q$, as the semantic constraints in the question are mainly defined by mentions of relations. Then, the rest of the tokens are fed to a pre-trained BERT model. [^4] https://huggingface.co/bert-base-uncased The last hidden states of the model are used as $\mathbf{Q}$. The output of this general question encoder is the general question embedding $\mathbf{q}\in\mathbb{R}^{d_{K}}$, where $d_{K}$ is the final embedding dimension. We implement the encoder as a bidirectional single-layer gated recurrent unit (GRU) network [34] and use the mean of the final GRU hidden states from two directions as $\mathbf{q}$.
The layerwise question encoder provides layerwise reference information about the given question for the KG encoding. In different layers, the reference information may have different foci. For example, given the question “who directed a movie starring Michael Keaton,” the reference information could have a focus on “starring” in the first layer and “directed a movie” in the second layer. This module is implemented as another bidirectional single-layer GRU network with the initial question encoding $\mathbf{Q}$ as input. The mean of its two final hidden states ($\mathbf{c}^{k}_{0}$ and $\mathbf{c}^{k}_{1}$) is used as the reference information $\mathbf{q}^{k}\in\mathbb{R}^{d_{k}}$ in the k-th layer. Since the model needs to track what has been focused on before, the final hidden states from the previous layer ($\mathbf{c}^{k-1}_{0}$ and $\mathbf{c}^{k-1}_{1}$) are used as initial hidden states of the current layer. As for the first layer, zero vectors are used as initial hidden states.
The attentional message propagation module computes messages that KG entities receive to update their embeddings in different layers. The input of this module in the $k$-th layer is the entity embedding matrix $\mathbf{E}^{k-1}\in\mathbb{R}^{|E|\times d_{k-1}}$ and the relation embedding matrix $\mathbf{R}^{k-1}\in\mathbb{R}^{|R|\times d_{k-1}}$ from the previous layer. $|E|$ and $|R|$ are the number of entities and the number of relations in the KG, respectively. Therefore, $\mathbf{E}^{k-1}$ and $\mathbf{R}^{k-1}$ contain one embedding vector for each entity and one for each relation, respectively. For the first layer, we initialize the relation embedding matrix $\mathbf{R}^{0}$ based on relation labels. Specifically, for each relation, we feed its label to the pre-trained BERT used for the question encoding and use a single-layer GRU network to encode the last hidden states of the BERT into one embedding vector. As for entities, the embeddings of topic entities are initialized by the final question embedding $\mathbf{q}$, while all other entities are initialized with a random vector that will be updated as parameters during the training of the whole GNN.
In this module, for each entity $e\in E$, we denote all its direct neighbors in the KG as $\mathcal{N}_{e}$. The messages that this entity receives and the aggregation of the messages are computed as follows: Regarding each neighboring entity $e_{i}\in\mathcal{N}_{e}$, if there is a relation $r_{j}\in R$ that links $e$ and $e_{i}$ in the KG regardless of the relation direction, i.e., $(e,r_{j},e_{i})\in T$ or $(e_{i},r_{j},e)\in T$, the entity $e$ receives a message $\mathbf{m}^{k}_{ij}$ in the k-th encoding layer:
$\mathbf{m}^{k}_{ij}=\operatorname{Tanh}\left(\mathbf{W}^{k}_{m}[\mathbf{e}^{k-1}_{i},\mathbf{r}^{k-1}_{j}]+\mathbf{b}^{k}_{m}\right),$ (1)
where $\operatorname{Tanh}(\cdot)$ denotes the hyperbolic tangent activation, $\mathbf{W}^{k}_{m}\in\mathbb{R}^{d_{k}\times 2d_{k-1}}$ and $\mathbf{b}^{k}_{m}\in\mathbb{R}^{d_{k}}$ are parameters of this module, and $[\mathbf{e}^{k-1}_{i},\mathbf{r}^{k-1}_{j}]$ denotes the concatenation of the embeddings of $e_{i}$ and $r_{j}$ from the previous layer. Then, the aggregation of all messages that $e$ receives is computed using an attention mechanism, where greater attention is paid to messages that are more relevant to the reference information $\mathbf{q}^{k}$. Specifically, for each message $\mathbf{m}^{k}_{ij}$, the attention score is computed as follows:
$s^{k}_{ij}=\mathbf{a}^{k}\operatorname{LeakyReLU}\left(\mathbf{W}^{k}_{a}[\mathbf{m}^{k}_{ij},\mathbf{q}^{k}]+\mathbf{b}^{k}_{a}\right),$ (2)
where $\mathbf{a}^{k}\in\mathbb{R}^{1\times d_{k}}$, $\mathbf{W}^{k}_{a}\in\mathbb{R}^{d_{k}\times 2d_{k}}$, $\mathbf{b}^{k}_{a}\in\mathbb{R}^{d_{k}}$ are parameters of the attention mechanism, and $[\mathbf{m}^{k}_{ij},\mathbf{q}^{k}]$ is the concatenation of $\mathbf{m}^{k}_{ij}$ and $\mathbf{q}^{k}$. Based on the Softmax function, the aggregated message for the entity $e$ is computed as follows:
$\mathbf{m}^{k}=\frac{1}{S^{k}}\sum_{e_{i}\in\mathcal{N}_{e}}\sum_{r_{j}\in\mathcal{R}_{e_{i}}}\exp\left(s^{k}_{ij}-s^{k}_{max}\right)\mathbf{m}^{k}_{ij},$ (3)
$\mathcal{R}_{e_{i}}=\{r_{j}\mid(e,r_{j},e_{i})\in T\lor(e_{i},r_{j},e)\in T\},$ (4)
$s^{k}_{max}=\max\left(\{s^{k}_{ij}\mid e_{i}\in\mathcal{N}_{e},r_{j}\in\mathcal{R}_{e_{i}}\}\right),$ (5)
$S^{k}=\sum_{e_{i}\in\mathcal{N}_{e}}\sum_{r_{j}\in\mathcal{R}_{e_{i}}}\exp\left(s^{k}_{ij}-s^{k}_{max}\right).$ (6)
In the above equations, $\mathcal{R}_{e_{i}}$ denotes the set of relations linking $e$ and $e_{i}$ in the KG.
The gated embedding update module includes an attentional gate that determines to what extent entity embeddings should be updated with aggregated messages computed by Eq. 3. This module takes the final question embedding $\mathbf{q}$ as a reference. Specifically, for the entity $e$ that receives $\mathbf{m}^{k}$ in the k-th layer, two attention scores $s^{k}_{m}$ and $s^{k}_{e}$ are respectively computed for $\mathbf{m}^{k}$ and the current entity embedding $\mathbf{e}^{k-1}$:
$s^{k}_{m}=\mathbf{a}^{k}_{u}\operatorname{LeakyReLU}\left(\mathbf{W}^{k}_{u}\mathbf{W}^{k}_{t1}[\mathbf{m}^{k},\mathbf{q}]+\mathbf{b}^{k}_{u}\right),$ (7)
$s^{k}_{e}=\mathbf{a}^{k}_{u}\operatorname{LeakyReLU}\left(\mathbf{W}^{k}_{u}\mathbf{W}^{k}_{t2}[\mathbf{e}^{k-1},\mathbf{q}]+\mathbf{b}^{k}_{u}\right),$ (8)
where $\mathbf{a}^{k}_{u}\in\mathbb{R}^{1\times d_{k}}$, $\mathbf{W}^{k}_{u}\in\mathbb{R}^{d_{k}\times 2d_{k}}$, $\mathbf{W}^{k}_{t1}\in\mathbb{R}^{2d_{k}\times(d_{k}+d_{K})}$, $\mathbf{W}^{k}_{t2}\in\mathbb{R}^{2d_{k}\times(d_{k-1}+d_{K})}$, $\mathbf{b}^{k}_{u}\in\mathbb{R}^{d_{k}}$ are parameters of this module. Please note that $\mathbf{W}^{k}_{t1}$ and $\mathbf{W}^{k}_{t2}$ are just for dimension conversion and should be fixed to identity matrices when $d_{k}=d_{k-1}=d_{K}$. Finally, the embedding of $e$ is updated as follows:
$\mathbf{e}^{k}=\frac{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)\mathbf{m}^{k}+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)\mathbf{e}^{k-1}}{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)},$ (9)
where $s^{\prime k}_{max}=\max(s^{k}_{m},s^{k}_{e})$.
We perform the above computations, i.e., Eqs. 1, 2, 3, 4, 5, 6, 7, 8 and 9, for each entity in the KG and obtain the updated entity embedding matrix $\mathbf{E}^{k}$. As for the relation embedding matrix, if the dimension of the k-th layer is changed, i.e., $d_{k}\neq d_{k-1}$, a linear transformation can be applied to $\mathbf{R}^{k-1}$. Otherwise, relation embeddings can be simply passed to the next layer.
Training and Inference. For a training question $q$ with the answer set $A_{q}\subseteq E$, we randomly sample pairs of answer and non-answer entities from the Cartesian product of $A_{q}$ and $E\setminus A_{q}$. The training objective is that, in the final embedding space, the Euclidean distance between $q$ and a correct answer in $A_{q}$ should be smaller than the Euclidean distance between $q$ and any non-answer entity sampled from $E\setminus A_{q}$ by a margin. Specifically, we define the loss of our GNN regarding $q$ as follows:
$loss(q)=\sum_{\left(e,e^{\prime}\right)\in A_{q}\times\left(E\setminus A_{q}\right)}\max\left(\|\mathbf{q},\mathbf{e}^{K}\|-\|\mathbf{q},\mathbf{e}^{\prime K}\|+\epsilon,0\right),$ (10)
where $e$ and $e^{\prime}$ respectively denote an answer and a non-answer entity, and $\epsilon\geq 0$ is the margin.
When the GNN is well-trained on all training questions, given a test question, we employ the GNN to encode the question and all entities into a joint embedding space. The top-N [^5] The value of N is a hyperparameter that can be set according to the performance of the GNN on the training/validation set. entities closest to the question are selected as candidate answers and passed to the next step.
IV-B Step-II: LM-based Explicit Reasoning
Figure 3: An overview of Step-II, assuming Tim Burton is extracted as a candidate answer regarding the example question and KG in Fig. 1.
Based on the example in Fig. 1, an overview of this step is depicted in Fig. 3. We first extract subgraphs from the given KG that consist of paths linking Tim Burton (candidate answer) to California and Michael Keaton (topic entities). Then, we rewrite the subgraphs into two natural language expressions and compare them with the given question to select the final reasoning subgraph according to their semantic similarity to the question.
[…]
Algorithm 1 Candidate reasoning subgraph extraction.
[…]
2:topic entity set $E_{q}$.
3:Candidate reasoning subgraph set $G_{q}$.
4:$G_{q}\leftarrow$ an empty set
5:for candidate answer $a_{c}\in A_{c}$ do
6:     if $a_{c}$ is connected to all topic entities in $E_{q}$ then
7:         $path\_sets\leftarrow$ an empty list
8:         for topic entity $e_{q}\in E_{q}$ do
[…]
10:         end for
[…]
12:              Add $\mathcal{G}_{c}$ to $G_{q}$
13:         end for
14:     end if
15:end for
[…]
Algorithm 2 Reasoning subgraph rewriting.
[…]
2:and topic entity set $E_{q}$.
3:Natural language expression set $\operatorname{NL}(G_{q})$.
4:$\operatorname{NL}(G_{q})\leftarrow$ an empty set
5:for subgraph $\mathcal{G}_{c}=(E_{c},R_{c},T_{c},\mathcal{P}_{c})\in G_{q}$ do
[…]
7:     for path $P$ in $\mathcal{P}_{c}$ do
8:         for triple $(e_{h},r,e_{t})$ in $P$ do
9:              if $r$ is an inversed relation then
[…]
11:              else
[…]
13:              end if
14:              if $e_{t}\in E_{q}$ then
[…]
16:              else
17:                  $nl\leftarrow nl+$ an entity that
18:              end if
19:         end for
20:         Remove $P$ from $\mathcal{P}_{c}$
21:         if $\mathcal{P}_{c}$ is not empty then
22:              $nl\leftarrow nl+$ and
23:         end if
24:     end for
25:     Add $nl$ to $\operatorname{NL}(G_{q})$
26:end for
Algorithm 3 The prediction of positive and negative natural language expressions.
[…]
3:natural language expressions.
4:$\operatorname{NL}_{p}(q)\leftarrow$ an empty set
5:$\operatorname{NL}_{n}(q)\leftarrow$ an empty set
6:$nl2ans\leftarrow$ an empty dictionary
7:$nl2vote\leftarrow$ an empty dictionary
8:$A_{c}\leftarrow$ candidate answer set of $q$ computed via Step-I
9:for entity $e\in A_{c}\cup A_{q}$ do
10:     $G_{q}\leftarrow$Longunderstack[l]the set of candidate reasoning subgraphs ex-
11:tracted with $e$ using Algorithm 1
12:     for subgraph $\mathcal{G}_{c}\in G_{q}$ do
13:         $Q(\mathcal{G}_{c})\leftarrow$Longunderstack[l]convert $\mathcal{G}_{c}$ into a query by turning $e$ in-
14:to a variable
15:         $A(\mathcal{G}_{c})\leftarrow$Longunderstack[l]a set of entities that can be obtained by
16:executing $Q(\mathcal{G}_{c})$ over $\mathcal{G}$
17:         $nl\leftarrow$Longunderstack[l]the natural language expression generated
18:based on $\mathcal{G}_{c}$ using Algorithm 2
19:         if $nl\notin keys(nl2ans)$ then
20:              $nl2ans[nl]\leftarrow$ an empty set
21:         end if
22:         for entity $e^{\prime}\in A(\mathcal{G}_{c})$ do
23:              Add $e^{\prime}$ to $nl2ans[nl]$
24:         end for
25:     end for
26:end for
27:for $nl\in$ $keys(nl2ans)$ do
28:     $upvote\leftarrow\left|nl2ans[nl]\cap A_{q}\right|$
29:     $downvote\leftarrow\left|nl2ans[nl]\setminus A_{q}\right|$
30:     if $upvote>0$ then
31:         $nl2vote[nl]\leftarrow upvote-downvote$
32:     end if
33:end for
34:$max\_vote\leftarrow$Longunderstack[l]maximum vote of any $nl\in keys(nl2vote)$
35:$min\_len\leftarrow$Longunderstack[l]minimum number of entity/relation ment-
36:ions in the expressions that have $max\_vote$
37:for $nl\in$ $keys(nl2vote)$ do
[…]
39:         Add $nl$ to $\operatorname{NL}_{p}(q)$
40:     else
41:         Add $nl$ to $\operatorname{NL}_{n}(q)$
42:     end if
43:end for
Semantic similarity comparison. In this step, we first use the Sentence Transformer [29] to encode the given question $q$ and natural language expressions in $\operatorname{NL}(G_{q})$ into embedding vectors. The output question embedding is denoted as $\mathbf{q}_{t}\in\mathbb{R}^{d_{t}}$, where $d_{t}$ is the output dimension of the Sentence Transformer. For an expression $nl\in\operatorname{NL}(G_{q})$, its output embedding is denoted as $\mathbf{nl}\in\mathbb{R}^{d_{t}}$. We select the optimal expression $nl^{\prime}$ that is closest to the question according to the output embeddings:
$nl^{\prime}=\operatorname*{arg\,max}_{nl\in\operatorname{NL}(G_{q})}\operatorname{sim}\left(\mathbf{nl},\mathbf{q}_{t}\right),$ (11)
where $\operatorname{sim}(\mathbf{nl},\mathbf{q}_{t})$ denotes the function that computes a similarity metric between $\mathbf{nl}$ and $\mathbf{q}_{t}$. Depending on which specific pre-trained model of the Sentence Transformer is used, the similarity metric could be dot product, cosine similarity, or the negative Euclidean distance. In our implementation, we adopt multi-qa-distilbert-cos-v1, [^7] https://huggingface.co/sentence-transformers/multi-qa-distilbert-cos-v1 which is a model pre-trained on the semantic search task [29] with more than 214M question-text answer pairs. The output dimension is 768, i.e., $d_{t}=768$, and $\operatorname{sim}\left(\mathbf{nl},\mathbf{q}_{t}\right)$ computes cosine similarity.
Finally, we return the subgraph in $G_{q}$ that corresponds to the optimal expression $nl^{\prime}$ as the final reasoning subgraph, and the candidate answer in this subgraph would be returned as the final answer. It is worth mentioning that, in practice, there could be several subgraphs corresponding to the same expression $nl^{\prime}$, as candidate answers and intermediate entities are not expressed when rewriting subgraphs in Algorithm 2. For example, besides Batman 1989, Beetlejuice is another movie directed by Tim Burton and starring Michael Keaton. For the question in Fig. 3, a reasoning subgraph for Beetlejuice would also be returned if it is contained in the KG. Therefore, all corresponding subgraphs and their matching KG entities could be considered as correct. However, if a more restricted set of answers (e.g., top-1 or top-10) is desired, the distances of the answers to the given question in the embedding space of Step-I can be used for ranking.
Sentence Transformer Fine-tuning. As aforementioned, the Sentence Transformer was pre-trained on the semantic search task that is different from our similarity comparison between questions and natural language expressions of subgraphs. To mitigate this gap, we fine-tune the Sentence Transformer based on the available weak supervision (i.e., question-final answer pairs).
[…]
As Algorithm 3 only predicts relatively positive and relatively negative expressions, the loss for fine-tuning the Sentence Transformer is accordingly defined as a triplet loss:
$l(q)=\sum_{\left(nl_{p},nl_{n}\right)\in\operatorname{NL}^{\prime}}\max\left(\operatorname{sim}\left(\mathbf{nl}_{n},\mathbf{q}_{t}\right)-\operatorname{sim}\left(\mathbf{nl}_{p},\mathbf{q}_{t}\right)+\epsilon^{\prime},0\right),$ (12)
where $\operatorname{NL}^{\prime}=\operatorname{NL}_{p}(q)\times\operatorname{NL}_{n}(q)$, $nl_{p}$ and $nl_{n}$ respectively denote a positive expression and a negative expression, $\operatorname{sim}(\cdot)$ is the function for similarity comparison from Eq. 11, and $\epsilon^{\prime}$ is a non-negative margin.
V Experiments
In this section, we evaluate the effectiveness and efficiency of the proposed model, examine the contributions of different model components, present a failure analysis, and demonstrate the explainability of generated reasoning subgraphs both qualitatively and quantitatively.
V-A Effectiveness Evaluation
TABLE I: The number of questions (in per split) and the average number of entities/relations/triples in question-specific subgraphs.
Dataset Question Set KG Subgraph (Average)
#Train #Valid #Test #Entity #Relation #Triple
WQSP 2,848 250 1,639 1,459 296 4,347
CWQ 27,639 3,519 3,531 1,473 292 4,751
PQ-2hop 1,527 190 191 1,056 13 1,211
PQ-3hop 4,159 519 520 1,836 13 2,839
PQL-2hop 1,276 159 159 5,034 363 4,247
PQL-3hop 825 103 103 6,505 411 5,597
Datasets. We used four benchmark datasets that are commonly adopted in the multi-hop QA task: WebQuestionsSP (WQSP) [35], ComplexWebQuestions (CWQ) [36], PathQuestion (PQ) [19], and PathQuestion-Large (PQL) [19]. [^8] Datasets available at https://github.com/ruijie-wang-uzh/GNN2R. WQSP and CWQ contain questions of varying complexity that require reasoning over up to two and four hops, respectively. PQL is a more challenging version of PQ with larger KGs and less training data. Both contain two subsets PQ(L)-2hop and PQ(L)-3hop, in which answering requires reasoning over two and three hops, respectively.
These QA datasets are all based on Freebase [37]. Following the baselines [13, 17, 23, 21], for each question in WQSP and CWQ, topic entities are identified using STAGG [38], and a subgraph consisting of relevant Freebase triples are extracted based on Personalized PageRank [39]. In PQ and PQL, questions are already annotated with topic entities. Also, medium-sized Freebase subgraphs are provided for each subset (i.e., PQ/PQL-2/3hop). Table I reports statistics of the adopted QA datasets, including the average number of entities/relations/triples in underlying KG subgraphs.
Hyperparameter Setting. We determined the following hyperparameters empirically via grid search. The number of GNN layers was always set to 3. The margin in Eq. 10 was set to 1.0 for WQSP, PQ-3hop, and both subsets of PQL, while 0.5 for CWQ and PQ-2hop. In the first step of GNN2R, the number of candidate answers to select was 10 for WQSP, 25 for CWQ, 20 for PQ-2hop, and 5 for PQ-3hop and both subsets of PQL. The margin in Eq. 12 was set to 0.1 for WQSP, CWQ, and PQ-3hop, 0.8 for PQ-2hop and PQL-2hop, and 0.05 for PQL-3hop.
Baselines. We compare with six baselines: EmbedKGQA [15], GraftNet [13], Relation Learning [17], TransferNet [21], NSM [23], and SR+NSM [40]. As aforementioned in Section II, EmbedKGQA, GraftNet, and Relation Learning are retrieval-based methods that learn KG representations and scoring functions to predict the likelihood of each entity being the answer to a given question. These methods are very similar to the first step of GNN2R, as we learn entity embeddings and use Euclidean distance for candidate answer selection. The other three baselines, i.e., TransferNet, NSM, and SR+NSM, are reasoning-based methods that can provide reasoning process over topological structures of KGs as explanations of returned answers. They are similar to the second step of GNN2R. It is worth mentioning that, in the weakly-supervised multi-hop QA task, these reasoning-based methods are state-of-the-art (SOTA) baselines.
TABLE II: Overall performance (in terms of Hits@1 and F1 score) on benchmark datasets. (best performance in bold, second best underlined, also percentage changes of our model with respect to the best baseline are shown at the bottom)
Method WQSP CWQ PQ-2hop PQL-2hop PQ-3hop PQL-3hop
Hits@1 F1 Hits@1 F1 Hits@1 F1 Hits@1 F1 Hits@1 F1 Hits@1 F1
EmbedKGQA [15] 66.6* 51.3 21.7 18.8 90.1 89.3 79.2 76.5 86.2 84.5 61.2 58.9
GraftNet [13] 68.7* 62.3* 36.8* 32.7* 90.6 89.3 87.1 74.3 93.1 92.7 84.0 65.1
Relation Learning [17] 72.9* 64.5* 35.1 31.3 99.5 98.4 98.1 97.7 59.0 53.9 85.4 85.7
TransferNet [21] 71.4* 66.1 48.6* 36.0 91.1 91.1 54.7 50.3 96.5 96.7 62.1 64.3
NSM [23] 74.3* 67.4* 48.8* 44.0* 94.2 93.7 74.2 75.0 97.1 96.2 67.0 68.3
SR+NSM [40] 69.5* 64.1* 50.2* 47.1* 93.7 93.5 76.7 75.7 96.7 96.7 68.0 66.3
GNN2R (Step-I) 68.9 46.4 48.8 43.3 96.9 95.5 99.4 80.9 94.8 88.8 91.3 84.5
GNN2R (Step-I + Step-II) 75.8 67.6 53.8 51.7 99.5 99.5 99.4 98.4 99.2 99.0 97.1 96.9
$\uparrow$ 2.0% $\uparrow$0.3$\uparrow$ 7.2% $\uparrow$ 9.8% 0.0% $\uparrow$ 1.1% $\uparrow$ 1.3% $\uparrow$0.7$\uparrow$ 2.2% $\uparrow$ 2.4% $\uparrow$ 13.7% $\uparrow$ 13.1%
Results marked with * are the best-reported results known to us from existing publications. Other results are obtained from our experiments using released source code and, when available, pre-trained models from the authors. In cases where pre-trained models are not available, we adopt the best model and training configurations reported in original papers and source code.
Overall Results. The overall performance of our model is reported in Table II. Following the baselines, we use Hits@1 and F1 score (both in percentage form) as evaluation metrics. Hits@1 refers to the proportion of test questions for which the top-1 answers returned are correct. Since there could be multiple correct answers for a given question, the F1 score is calculated based on the precision and recall of all returned answers. In the row GNN2R (Step-I), we report the results with only the first step of our model. Specifically, for a test question, we return its nearest-neighboring entity in the computed embedding space as the top-1 answer. For calculating the F1 score, other entities within a threshold of distance [^9] This distance is calculated by multiplying the shortest distance by a pre-defined threshold. are returned as all possible answers. In the row GNN2R (Step-I + Step-II), we report the results of the complete model. As introduced in Section IV-B, final answers (entities) can be obtained based on reasoning subgraphs. Similarly, the entity that has the shortest distance in Step-I is returned as the top-1 answer, while other entities within a threshold of distance form the complete answer set.
The following can be observed from Table II:
1) GNN2R achieves superior performance on all benchmark datasets in comparison with all six baselines. Especially on CWQ and PQL-3hop, which are the two most challenging datasets in our experiments regarding question complexity and the size of KGs, significant improvements (7.2%/9.8% on CWQ and 13.7%/13.1% on PQL-3hop with respect to Hits@1/F1) have been achieved in comparison to previous best-performing baselines. Also, it is worth mentioning that PQ and PQL are nearly solved by GNN2R (cf. Table VI and Section V-D). To the best of our knowledge, we are the first and, thus far, the only one to achieve this. [^10] A public leaderboard is available at https://github.com/KGQA/leaderboard/blob/gh-pages/freebase/path_question.md. These results demonstrate the overall effectiveness of our method.
2) GNN2R remains highly competitive even with only the first step (Step-I). For example, on CWQ, Step-I achieves a similar performance as the previous SOTA method NSM and outperforms four other baselines (EmbedKGQA, GraftNet, Relation Learning, and TransferNet). On PQL-2hop and PQL-3hop, Step-I is ranked first among all baselines regarding Hits@1. Also, if considering F1, Step-I is ranked second on PQ-2hop, PQL-2hop, and PQL-3hop. These results demonstrate that the proposed GNN and GNN-based coarse reasoning in Step-I are highly effective.
3) The performance of GNN2R can be substantially improved with the second step (Step-II). Specifically, on the six benchmark datasets, average improvements of 5.7% and 19.5% regarding Hits@1 and F1 are achieved when Step-II is performed based on Step-I. These results demonstrate the overall effectiveness of Step-II and indicate that the proposed algorithms are effective in fine-tuning the LM for selecting reasoning subgraphs and, consequently, correct answers. Moreover, it is noteworthy that the improvement in F1 is more substantial than that in Hits@1 (19.5% v.s. 5.7%). This is in line with our intuition that Step-II can effectively mitigate the bottleneck of Step-I in selecting multiple correct answers from a continuous search space, as the retrieved reasoning subgraph can prune all candidate answers that do not satisfy the semantic constraints of given questions.
4) GNN2R demonstrates superior robustness to all baselines when the underlying KG becomes bigger and the training data becomes smaller. From PQ-2hop to PQL-2hop, the average number of underlying KG triples increases by 250.7%, and the number of total training questions decreases by 16.4%. Consequently, the performance of the six baselines drops by 16.1% and 19.3% with respect to Hits@1 and F1 on average. In contrast, GNN2R only drops by 0.1% (Hits@1) and 1.1% (F1). From PQ-3hop to PQL-3hop, the average number of triples increases by 97.1%, and the number of training questions decreases by 80.2%. In this case, the performance of the six baselines drops by 15.1% (Hits@1) and 15.8% (F1) on average. In contrast, GNN2R only drops by 2.1% (Hits@1) and 2.1% (F1).
TABLE III: The average run-time (ms) of all major phases of GNN2R in the process of answering a question. (mean $\mu$ and standard deviation $\sigma$ of five runs)
Step Phases $\mu$ $\sigma$
Step-I: GNN-based Question & KG Preprocessing 22.1 0.2
Coarse Reasoning GNN Encoding 5.7 0.0
Candidate Answer Selection 3.4 0.1
Step-II: LM-based Reasoning Subgraph Extraction 160.1 1.2
Explicit Reasoning Reasoning Subgraph Rewriting 0.3 0.0
Semantic Similarity Comparison 28.7 0.1
Total (including data loading, etc.) 222.1 1.3
TABLE IV: Comparison of the average run-time (ms) for answering a question. (mean $\mu$ and standard deviation $\sigma$ of five runs)
GNN2R GNN2R-fast NSM TransferNet
$\mu$ 222.1 200.9 241.4 223.5
$\sigma$ 1.3 0.9 4.7 18.4
V-B Efficiency Evaluation
In this section, we aim to evaluate the efficiency of GNN2R in answering questions. As WQSP is collected from Google Suggest API [41] with human evaluations on Amazon Mechanical Turk, [^11] https://www.mturk.com/ we consider it a good representation of real-world questions and use it in this evaluation. The average run-time of GNN2R, along with the time costs associated with all major phases of the answering process, is reported in Table III. On average, GNN2R can answer a test question within 222.1ms. Considering that 100ms is typically seen as instantaneous and less than 1s as non-interrupting in interaction design, [^12] https://www.nngroup.com/articles/powers-of-10-time-scales-in-ux/ and https://www.nngroup.com/articles/response-times-3-important-limits/ GNN2R is highly efficient and applicable in real-world scenarios.
We also compare the efficiency of GNN2R with two previous SOTA methods (NSM and TransferNet). [^13] SR+NSM aims to provide enhanced input subgraphs for NSM. It shares the same reasoning process as NSM and, therefore, is not distinguished here. Table IV reports the average run-time of GNN2R, GNN2R-fast, NSM, and TransferNet on the test set of WQSP. GNN2R-fast refers to a faster version of GNN2R, which only considers the shortest paths between topic entities and candidate answers when extracting candidate reasoning subgraphs in Algorithm 1. The Hits@1 of GNN2R-fast on WQSP slightly drops to 75.4, which is still higher than NSM, TransferNet, and SR+NSM according to Table II. It can be observed that GNN2R is substantially more efficient than NSM (with an 8.0% relative time reduction) and slightly more efficient than TransferNet. GNN2R-fast is substantially more efficient than both NSM and TransferNet, with 16.8% and 10.1% relative time reductions, respectively. Also, the run-times for GNN2R and GNN2R-fast seem to be more stable across different runs, as suggested by the smaller standard deviations.
The above evaluations were conducted on a Linux server with two Intel Xeon Gold 6230 CPUs (2.10GHz) and one NVIDIA GeForce RTX 2080 Ti GPU being used. For NSM and TransferNet, the code released by original authors is used. [^14] https://github.com/RichardHGL/WSDM2021_NSM and https://github.com/shijx12/TransferNet Time for generating answer explanations is included for all models. [^15] Please refer to Section V-E for how the baselines generate explanations.
TABLE V: The performance of GNN2R in successively degraded scenarios. (Hits@1 %)
Scenario WQSP CWQ
Complete Step-I and Step-II 75.8 53.8
Step-I and Step-II without Fine-tuning 60.1 39.1
Only Step-I 68.9 48.8
Only Step-I with Degraded GNN 62.8 47.5
Only Step-I with Further Degraded GNN 56.4 38.6
TABLE VI: Statistics of failed questions and their corresponding failure phase.
Failure Phase WQSP CWQ PQ-2hop PQL-2hop PQ-3hop PQL-3hop
Errors in entity linking and the preparation of KG subgraphs 88 (22.2%) 920 (56.4%) 0 (0.0%) 0 (0.0%) 0 (0.0%) 0 (0.0%)
Errors in the GNN-based candidate answer retrieval 97 (24.5%) 63 (3.9%) 0 (0.0%) 0 (0.0%) 1 (25.0%) 1 (33.3%)
Errors in the LM-based reasoning subgraph retrieval 211 (53.3%) 648 (39.7%) 1 (100%) 1 (100%) 3 (75.0%) 2 (66.7%)
Total number of failed questions 396 1,631 1 1 4 3
V-C Ablation Study
To examine the contributions of various components proposed in GNN2R, we conducted an ablation study on WQSP and CWQ. The results are reported in Table V. Following the complete version of GNN2R in the first row of the table, there are four successively degraded scenarios:
1) In the second row, we examine the performance of the model without fine-tuning the LM. On WQSP and CWQ, the Hits@1 drops by 20.7% and 27.3%, respectively. This demonstrates that Algorithm 3 can indeed generate high-quality positive and negative expressions for the fine-tuning. Also, the gap between the pre-trained LM and our specific task in Step-II can be effectively mitigated.
2) The third row is the result of only Step-I. The contribution of Step-II can be shown by comparing it with the first row, which has also been analyzed in Section V-A. In addition, we can observe that the result in this row is substantially better in comparison with the second row, where Step-II is performed without fine-tuning the LM. This again shows the necessity and effectiveness of the proposed algorithms for fine-tuning.
3) In the fourth row, we further remove the gated embedding update module (cf. the blue box on the lower side in Fig. 2). In this case, entity embeddings are directly the aggregation of received messages, which is a common setup of GNNs. Compared with the above scenario, the performance drops by 8.9% on WQSP and 2.7% on CWQ, which demonstrates the contribution of our question-dependent embedding update.
4) In the last row, we further remove the layerwise question encoder (cf. the yellow box in Fig. 2) and the attentional message propagation module (cf. the blue box on the upper side in Fig. 2). In this case, the GNN degrades to a basic GNN structure similar to GCN [42]. We can observe a further performance drop of 10.2% on WQSP and 18.7% on CWQ, which demonstrates the contribution of the question-dependent message propagation.
V-D Failure Analysis
We categorize failed questions based on the processing phase in which the error occurs: a) errors in entity linking and the preparation of KG subgraphs, b) errors in the GNN-based candidate answer retrieval, c) errors in the LM-based reasoning subgraph retrieval. Table VI reports statistics of failed benchmark questions. The following can be observed:
1) For WQSP, most errors (53.3%) occur in the reasoning subgraph retrieval phase. We believe the reason is two-fold: First, WQSP only provides 2,848 training questions that are insufficient for fine-tuning the Sentence Transformer, which was pre-trained on 215M question-answer pairs for a very different task. [^16] More details available at https://huggingface.co/sentence-transformers/multi-qa-distilbert-cos-v1. Second, there are no ground-truth annotations of reasoning subgraphs available in our task. The fine-tuning is based on the weak supervision inferred from Algorithm 3, which inevitably contains errors and noise.
2) For CWQ, most errors (56.4%) occur in the pre-processing phase where external tools are used. This is mainly due to the language complexity of the questions in CWQ. Also, it shows that there is substantial room for our model to improve with better pre-processing. Furthermore, in comparison to WQSP, the ratio of errors in subgraph retrieval drops by 25.5%. This increased performance of the LM could be mainly driven by the bigger training set of CWQ (27,639 questions).
3) For CWQ, only 3.9% errors occur in the GNN-based candidate answer retrieval, which shows the effectiveness of the proposed GNN considering the complexity of CWQ questions. For WQSP, a higher ratio of errors (24.5%) occurs in this phase, which could be due to the lack of training questions again.
4) For PQ and PQL, only 9 out of 973 total test questions are failed by our model. There are 2 and 7 failures in the GNN-based phase and LM-based phase, respectively. We could not observe any clear pattern from them and believe they are corner cases for the trained GNN and LM.
V-E Explainability Evaluation
In this section, we present a qualitative case study and a quantitative evaluation to demonstrate the explainability of generated reasoning subgraphs.
TABLE VII: Top-5 answers returned by Step-I for the example question.
Example Question Candidate Answer Distance
what state is saint louis university in? Missouri 6.08
St. Louis 6.78
United States of America 7.55
Pius XII Memorial Library 8.00
Saint Louis University 8.06
Positive Case. We present a positive case from WQSP: “what state is saint louis university in,” which has a topic entity Saint Louis University. In GNN2R Step-I, we use the GNN to encode the question and KG entities in a joint embedding space, where 10 nearest neighbors of the question are selected as candidate answers. The top-5 and their distances to the question are reported in Table VII. Then, in Step-II, 136 candidate reasoning subgraphs are extracted. For Missouri, the extracted subgraphs include: {(Missouri, administrative parent, United States of America), (United States of America, containedby-1, Saint Louis University)} and {(Missouri, state-1, St. Louis), (St. Louis, containedby-1, Saint Louis University)}. They are respectively rewritten into “what has the administrative parent an entity that is the containedby of Saint Louis University” and “what is the state of an entity that is the containedby of Saint Louis University.” [^17] Please note that the natural language expressions are only used as input for the LM and do not need to be grammatically correct. The semantic similarities computed by the fine-tuned LM for the two expressions are 0.20 and 0.66, respectively, which are consistent with human intuition. Finally, {(Missouri, state-1, St. Louis), (St. Louis, containedby-1, Saint Louis University)} is returned as the final reasoning subgraph, which explains the answer Missouri.
Negative Cases. We present a negative case for each of the failure phases in Table VI:
1) what are major exports of the usa? The answers, e.g., Automotive Industry and Food Manufacturing, are not included in the extracted Freebase subgraph.
2) what did st augustine do? In Step-I, the nearest neighbors of the question in the embedding space are mainly books related to the topic entity Augustine of Hippo, e.g., The Confessions of St. Augustine and The Enchiridion Manual. The expected answers, e.g., Writer and Physician, are not returned by Step-I.
3) what highschool did harper lee go to ? In Step-I, the correct answer Monroe County High School is successfully ranked the $5^{th}$ among 1,735 entities in the subgraph of this question. However, the final reasoning subgraph is {(Huntingdon College, students_graduates, m.0lwxmy1181818m.0lwxmy1 is a compound value type entity.), (m.0lwxmy1, education-1, Harper Lee)}, which returns the college instead of the requested high school of Harper Lee as the final answer.
TABLE VIII: Evaluation of explanations generated by GNN2R and other reasoning-based baselines.
PQ-2hop PQL-2hop PQ-3hop PQL-3hop
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
NSM 0.14 0.28 0.18 0.10 0.16 0.12 0.13 0.23 0.17 0.13 0.14 0.14
SR+NSM 0.13 0.27 0.18 0.08 0.11 0.09 0.15 0.25 0.19 0.12 0.13 0.12
TransferNet 0.89 0.89 0.89 0.39 0.45 0.42 0.84 0.94 0.89 0.50 0.60 0.55
GNN2R 0.97 0.97 0.97 0.84 0.69 0.76 0.91 0.89 0.90 0.89 0.63 0.74
$\uparrow$ 9.0% $\uparrow$ 9.0% $\uparrow$ 9.0% $\uparrow$ 115.4% $\uparrow$ 53.3% $\uparrow$ 81.0% $\uparrow$ 8.3% $\downarrow$ 5.3% $\uparrow$ 1.1% $\uparrow$ 78.0% $\uparrow$ 5.0% $\uparrow$ 34.5%
Explainability Comparison. We compare our reasoning subgraphs with explanations that the SOTA reasoning-based methods (i.e., NSM, SR+NSM, and TransferNet) can provide.
NSM and SR+NSM compute step-wise entity distributions during reasoning over KGs. At the initial state, topic entities are assigned high probabilities (1 divided by the number of topic entities). Then, entities with high probabilities during intermediate/final steps are assumed to be intermediate/answer entities in reasoning chains, which can be used as explanations. The main drawback is that NSM and SR+NSM cannot examine if the reasoning chains are semantically consistent with given questions. For example, given the question from WQSP “who is gimli’s father in the hobbit,” the returned answer of the authors’ pre-trained NSM is correct: Gloin. However, the inferred most-likely reasoning chain is {(Gloin, species, Dwarf), (Dwarf, characters of this species, Gimli)}, which obviously does not match the given question. In GNN2R, as we employ an LM to compare the semantics of reasoning subgraphs and given questions, this issue is alleviated, and the correct explanation can be returned: {(Gloin, parents-1, Gimli)}.
TransferNet also propagates probabilities over KG topological structures to infer answers. At the initial state, topic entities are initialized with probabilities of 1. During reasoning, step-wise transfer scores are computed for relations, which are used to propagate probabilities from topic entities to final answers. The authors propose to extract relations with transfer scores higher than a threshold at each step to explain the reasoning process. However, this threshold is difficult to specify, and there are usually too many relations returned with very close scores. For example, in the above-mentioned question “what state is saint louis university in,” TransferNet performs a two-step reasoning. With the threshold of 0.9 specified in the released implementation, 204 and 225 relations are selected, respectively. In the first step, top selected relations (with computed scores) include headquarters (0.999), institution (0.997), location (0.997), and citytown (0.997). In the second step, top selected relations include location (1.000), institution (1.000), and spouse_s_reverse (1.000). It is difficult and tedious for users to examine these relations. In comparison, our reasoning subgraph is straightforward and concise, as presented above.
Explanation Evaluation. We use ground-truth annotations of reasoning chains in PQ and PQL to further examine the quality of our generated reasoning subgraphs. A comparison with recent reasoning-based methods, including NSM, SR+NSM, and TransferNet, is reported in Table VIII. Specifically, we check if the triples within our reasoning subgraphs or the explanations of the baselines are present in the ground-truth reasoning chains and calculate precision, recall, and F1 on this triple level. It can be observed that, in general, the quality of our generated reasoning subgraphs is substantially higher than that of existing reasoning-based methods. In comparison to the best-performing baseline, i.e., TransferNet, average improvements of 52.7% (precision), 15.5% (recall), and 31.4% (F1) are achieved on the four datasets. Also, on the datasets with less training data and bigger KGs, i.e., PQL-2hop and PQL-3hop, the average improvements of GNN2R become even larger: 96.7% (precision), 29.2% (recall), and 57.8% (F1). NSM and SR+NSM are not comparable to TransferNet and our model regarding the quality of generated explanations. The main reason is that NSM and SR+NSM do not examine if their reasoning chains are semantically consistent with given questions. As for TransferNet, the major weakness is the precision of explanations. As analyzed above, TransferNet tends to return redundant relations and, consequently, triples in final explanations. This is also the reason that it performed better than GNN2R on PQ-3hop regarding recall. However, a higher recall does not necessarily signify a better quality of explanation, especially given its lower precision.
VI Conclusion
In this paper, we focus on the problem of rationale-providing QA over KGs and propose a novel model called GNN2R to tackle two critical challenges in this task: weak supervision and the high-efficiency requirement. GNN2R can retrieve both final answers and reasoning subgraphs as explanations, providing a human-readable rationale for the answers. The model answers a given question in two steps: In Step-I GNN-based coarse reasoning, a novel GNN architecture is proposed to represent the given question and KG entities in a joint embedding space, where the question is close to its answers. This step can efficiently prune the search space of candidate answers and reasoning subgraphs. In Step-II LM-based explicit reasoning, simple and effective algorithms are proposed to extract candidate reasoning subgraphs and fine-tune an LM that selects final reasoning subgraphs via semantic similarity comparison. Extensive experiments have been conducted on commonly adopted benchmark datasets to evaluate the proposed model and its components, analyze failed cases, and examine the generated reasoning subgraphs. According to the experimental results, GNN2R outperforms recent SOTA methods that are applicable to this task in terms of effectiveness, efficiency, and quality of generated explanations. In future work, we aim to provide users with more natural and intuitive explanations and examine them in real-world user testing. This can be achieved by generating text summaries of answers and reasoning subgraphs using generative language models.
Acknowledgments
This work has been partially supported by the University Research Priority Program “Dynamics of Healthy Aging” at the University of Zurich and the Swiss National Science Foundation through project MediaGraph (contract no. 202125). Michael Cochez was in part supported by the Horizon Europe Graphmassivizer project under grant agreement 101093202.

### 20 Jan 2024 v3

@@
 Index Terms:
-Knowledge Graphs, Question Answering, Graph Neural Networks, Language Models
+Knowledge Graph, Question Answering, Graph Neural Network, Language Model
 I Introduction
-In recent years, knowledge graphs (KGs) have been prominently used as knowledge sources supporting multi-hop question-answering (QA) [1].
+In recent years, knowledge graphs (KGs) have been prominently used to organize and provide knowledge for question answering (QA) tasks [1, 2, 3, 4, 5].
 An example is given in Fig.
@@
 This seems to be a successful application of KG-based QA.
-However, it is difficult to verify and understand the answer for normal users, given the complexity of navigating question-relevant parts within a large KG, not to mention the technical expertise required for such navigation, e.g., SPARQL [2] and understanding of the schema of the KG [3].
+However, it is difficult to verify and understand the answer for normal users, given the complexity of navigating question-relevant parts within a large KG, not to mention the technical expertise required for such navigation, e.g., SPARQL [6] and understanding of the schema of the KG.
 We propose to provide users with a reasoning subgraph consisting of triples from the KG that elucidate how the answer satisfies the semantic constraints specified in the question.
@@
 II Related Work
-We introduce three types of current mainstream QA methods and discuss the potential application of large language models.
-Query construction-based methods [4, 5, 6, 7, 8, 9, 10] try to translate given questions into logical queries, e.g., SPARQL [2] and symbolic programs [11], that can be executed over KGs to obtain final answers.
-They have been studied for a long time and have achieved competitive performance.
+In this section, we introduce three types of current mainstream KG-based QA methods and discuss the application of large language models (LLMs) in this domain.
+Query construction-based methods [7, 8, 9, 10] translate given questions into logical queries, e.g., SPARQL [6] and symbolic programs [11], that can be executed over KGs to obtain answers.
+This type of method has been extensively studied since the inception of KG-based QA.
 Also, by instantiating the constructed logical queries, these methods can be potentially used to generate reasoning subgraphs.
-However, recent methods of this type, especially case-based [6], transfer learning-based [7], and supervised learning-based [10], heavily rely on annotations of logical queries, which makes them inapplicable in low-resource scenarios with only weak supervision available, including the task of this paper, where only final answers are annotated for training.
-Retrieval-based methods [12, 13, 14, 15, 16, 17] compute contextual KG representations and leverage distance computations or learned neural classifiers to retrieve answers by the parametric ranking or classification of entities.
+However, recent SOTA methods of this type, including case-based [7], transfer learning-based [8], and supervised learning-based [10], heavily rely on existing corpora or annotations of logical queries, making them inapplicable to the task of this paper, where only weak supervision (i.e., annotations of final answers) is available for training.
+Retrieval-based methods [12, 13, 14, 15] directly retrieve answers from KGs via parametric ranking or classification of KG entities based on their contextual representations, e.g., KG embeddings.
 For example, Sun et al.
 [13] propose to use a graph neural network [12] to compute contextual representations of KG entities and use these embeddings to classify entities into answers and non-answer entities.
-These methods are highly efficient.
-However, as these methods do not explicitly reason over triples, they do not offer explanations.
-In the first step, our model selects candidate answers from computed embedding spaces, which aligns with this category.
-Reasoning-based methods [18, 19, 20, 21, 22, 23] are potentially capable of extracting reasoning subgraphs from KGs.
-A branch [18, 20] of these methods performs explicit explorations in KGs and generates reasoning chains for given questions.
+These methods are usually highly efficient, as they employ end-to-end numerical computations without explicit reasoning over KGs.
+Consequently, they lack interpretability and cannot provide explicit answer explanations.
+The first step of our model selects candidate answers from computed embedding spaces, which is in line with this type of method.
+However, the above weaknesses are effectively mitigated by the second step.
+Reasoning-based methods [16, 17, 18, 19, 20, 21] can be divided into two branches.
+A branch [16, 18] performs explicit explorations in KGs and generates reasoning chains for given questions.
 For example, Qiu et al.
 [18] use a reinforcement learning agent that explores paths in the KG starting from topic entities mentioned in given questions.
-Another branch [21, 22] performs fuzzy reasoning by computing confidence scores or numerical labels for entities and triples in KGs.
+Another branch [19, 20] performs fuzzy reasoning by computing confidence scores or numerical labels for KG entities and triples.
 For example, Shi et al.
@@
 Furthermore, most of these methods can only provide probable reasoning chains with likelihoods that are difficult to comprehend.
-The application of large language models (LLMs), in KG-based QA has recently attracted wide attention [24, 25, 26].
+The application of LLMs in KG-based QA has recently attracted wide attention [22, 23, 24].
 For example, Tan et al.
-[25] find that LLMs have the potential to outperform traditional QA methods, which do not rely on LLMs, in a relaxed setup, where a question is deemed correctly answered if the response from LLMs mentions any valid answer.
+[23] find that LLMs have the potential to outperform traditional QA methods that do not rely on LLMs in a relaxed setup, where a question is deemed correctly answered if the response from LLMs mentions any valid answer.
 However, this study also highlights a crucial drawback of LLMs—extracting answers from their textual response of varying length is difficult, especially considering that the answers may be expressed as aliases, acronyms, synonyms, or even in a different language.
 Also, given the implicit nature of the knowledge in LLMs and the common hallucination issue [25], it is challenging to verify the validity and completeness of answers from LLMs.
-In contrast, traditional QA methods can provide explicit answers in an unambiguous and uniform manner, and our method is able to extract the rationale behind answers from KGs.
+In contrast, traditional QA methods can provide explicit answers in an unambiguous and uniform manner.
+Also, our method is able to extract the rationale behind answers from KGs.
 Furthermore, Sun et al.
-[26] show that even the current state-of-the-art LLMs, including ChatGPT [^2] https://platform.openai.com/docs/api-reference and LLaMA (33B) [28], still have difficulty answering simple (one-hop) factual questions generated from templates, which is often considered a solved task with traditional QA methods.
+[24] show that even the current SOTA LLMs, including ChatGPT [^2] https://platform.openai.com/docs/api-reference and LLaMA (33B) [26], still have difficulty answering simple (one-hop) factual questions generated from templates, which is often considered a solved task with traditional QA methods.
 Therefore, in the era of LLMs, the contributions of this paper, alongside other methods without relying on LLMs, still hold practical significance.
@@
 Finally, the embedding of $e$ is updated as follows:
-$\mathbf{e}^{k}=\frac{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)\mathbf{m}^{k}+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)\mathbf{e}^{k-1}}{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)},$ (9)
-where $s^{\prime k}_{max}=\max(s^{k}_{m},s^{k}_{e})$.
+$\mathbf{e}^{k}=\frac{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)\mathbf{m}^{k}+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)\mathbf{W}^{k}_{t3}\mathbf{e}^{k-1}}{\exp\left(s^{k}_{m}-s^{\prime k}_{max}\right)+\exp\left(s^{k}_{e}-s^{\prime k}_{max}\right)},$ (9)
+where $s^{\prime k}_{max}=\max(s^{k}_{m},s^{k}_{e})$ and $\mathbf{W}^{k}_{t3}\in\mathbb{R}^{d_{k}\times d_{k-1}}$. Also, $\mathbf{W}^{k}_{t3}$ should be an identity matrix if $d_{k}=d_{k-1}$.
 We perform the above computations, i.e., Eqs.
@@
 26:end for
-Algorithm 3 The prediction of positive and negative natural language expressions.
-[…]
-[…]
-3:natural language expressions.
-4:$\operatorname{NL}_{p}(q)\leftarrow$ an empty set
-5:$\operatorname{NL}_{n}(q)\leftarrow$ an empty set
-6:$nl2ans\leftarrow$ an empty dictionary
-7:$nl2vote\leftarrow$ an empty dictionary
-8:$A_{c}\leftarrow$ candidate answer set of $q$ computed via Step-I
-9:for entity $e\in A_{c}\cup A_{q}$ do
-10:     $G_{q}\leftarrow$Longunderstack[l]the set of candidate reasoning subgraphs ex-
-11:tracted with $e$ using Algorithm 1
-12:     for subgraph $\mathcal{G}_{c}\in G_{q}$ do
-13:         $Q(\mathcal{G}_{c})\leftarrow$Longunderstack[l]convert $\mathcal{G}_{c}$ into a query by turning $e$ in-
-14:to a variable
-15:         $A(\mathcal{G}_{c})\leftarrow$Longunderstack[l]a set of entities that can be obtained by
-16:executing $Q(\mathcal{G}_{c})$ over $\mathcal{G}$
-17:         $nl\leftarrow$Longunderstack[l]the natural language expression generated
-18:based on $\mathcal{G}_{c}$ using Algorithm 2
-19:         if $nl\notin keys(nl2ans)$ then
-20:              $nl2ans[nl]\leftarrow$ an empty set
-21:         end if
-22:         for entity $e^{\prime}\in A(\mathcal{G}_{c})$ do
-23:              Add $e^{\prime}$ to $nl2ans[nl]$
-24:         end for
-25:     end for
-26:end for
-27:for $nl\in$ $keys(nl2ans)$ do
-28:     $upvote\leftarrow\left|nl2ans[nl]\cap A_{q}\right|$
-29:     $downvote\leftarrow\left|nl2ans[nl]\setminus A_{q}\right|$
-30:     if $upvote>0$ then
-31:         $nl2vote[nl]\leftarrow upvote-downvote$
-32:     end if
-33:end for
-34:$max\_vote\leftarrow$Longunderstack[l]maximum vote of any $nl\in keys(nl2vote)$
-35:$min\_len\leftarrow$Longunderstack[l]minimum number of entity/relation ment-
-36:ions in the expressions that have $max\_vote$
-37:for $nl\in$ $keys(nl2vote)$ do
-[…]
-39:         Add $nl$ to $\operatorname{NL}_{p}(q)$
-40:     else
-41:         Add $nl$ to $\operatorname{NL}_{n}(q)$
-42:     end if
-43:end for
 Semantic similarity comparison.
@@
 However, if a more restricted set of answers (e.g., top-1 or top-10) is desired, the distances of the answers to the given question in the embedding space of Step-I can be used for ranking.
+Algorithm 3 The prediction of positive and negative natural language expressions.
+[…]
+2:answer set $A_{q}$, computed results of GNN2R Step-I.
+[…]
+4:$\operatorname{NL}_{n}(q)$ of natural language expressions.
+5:$\operatorname{NL}_{p}(q)\leftarrow$ an empty set
+6:$\operatorname{NL}_{n}(q)\leftarrow$ an empty set
+7:$nl2ans\leftarrow$ an empty dictionary
+8:$nl2vote\leftarrow$ an empty dictionary
+9:$A_{c}\leftarrow$ candidate answer set of $q$ computed via Step-I
+10:for entity $e\in A_{c}\cup A_{q}$ do
+11:     $G_{q}\leftarrow$Longunderstack[l]the set of candidate reasoning subgraphs ex-
+12:tracted with $e$ using Algorithm 1
+13:     for subgraph $\mathcal{G}_{c}\in G_{q}$ do
+14:         $Q(\mathcal{G}_{c})\leftarrow$Longunderstack[l]convert $\mathcal{G}_{c}$ into a query by turning $e$ in-
+15:to a variable
+16:         $A(\mathcal{G}_{c})\leftarrow$Longunderstack[l]a set of entities that can be obtained by
+17:executing $Q(\mathcal{G}_{c})$ over $\mathcal{G}$
+18:         $nl\leftarrow$Longunderstack[l]the natural language expression generated
+19:based on $\mathcal{G}_{c}$ using Algorithm 2
+20:         if $nl\notin keys(nl2ans)$ then
+21:              $nl2ans[nl]\leftarrow$ an empty set
+22:         end if
+23:         for entity $e^{\prime}\in A(\mathcal{G}_{c})$ do
+24:              Add $e^{\prime}$ to $nl2ans[nl]$
+25:         end for
+26:     end for
+27:end for
+28:for $nl\in$ $keys(nl2ans)$ do
+29:     $upvote\leftarrow\left|nl2ans[nl]\cap A_{q}\right|$
+30:     $downvote\leftarrow\left|nl2ans[nl]\setminus A_{q}\right|$
+31:     $nl2vote[nl]\leftarrow upvote-downvote$
+32:end for
+33:$max\_vote\leftarrow$Longunderstack[l]maximum vote of any $nl\in keys(nl2vote)$
+34:$min\_len\leftarrow$Longunderstack[l]minimum number of entity/relation ment-
+35:ions in the expressions that have $max\_vote$
+36:for $nl\in$ $keys(nl2vote)$ do
+[…]
+38:         Add $nl$ to $\operatorname{NL}_{p}(q)$
+39:     else
+40:         Add $nl$ to $\operatorname{NL}_{n}(q)$
+41:     end if
+42:end for
 Sentence Transformer Fine-tuning.
@@
 A comparison with recent reasoning-based methods, including NSM, SR+NSM, and TransferNet, is reported in Table VIII.
-Specifically, we check if the triples within our reasoning subgraphs or the explanations of the baselines are present in the ground-truth reasoning chains and calculate precision, recall, and F1 on this triple level.
-It can be observed that, in general, the quality of our generated reasoning subgraphs is substantially higher than that of existing reasoning-based methods.
+Specifically, we check if the triples within our reasoning subgraphs or the explanations of the baselines are present in the ground-truth reasoning chains and calculate precision, recall, and F1 on the triple level.
+The following can be observed:
+1) It can be observed that, except for a performance drop on PQ-3hop regarding recall, the quality of our generated reasoning subgraphs is substantially higher than that of existing reasoning-based methods.
 In comparison to the best-performing baseline, i.e., TransferNet, average improvements of 52.7% (precision), 15.5% (recall), and 31.4% (F1) are achieved on the four datasets.
 Also, on the datasets with less training data and bigger KGs, i.e., PQL-2hop and PQL-3hop, the average improvements of GNN2R become even larger: 96.7% (precision), 29.2% (recall), and 57.8% (F1).
-NSM and SR+NSM are not comparable to TransferNet and our model regarding the quality of generated explanations.
+This demonstrates the strength of GNN2R in handling large search spaces with limited weak supervision.
+2) NSM and SR+NSM are not comparable to TransferNet and our model regarding the quality of generated explanations.
 The main reason is that NSM and SR+NSM do not examine if their reasoning chains are semantically consistent with given questions.
+As illustrated by the above example “who is gimli’s father in the hobbit,” the reasoning chains of NSM and SR+NSM are often incorrect, despite yielding correct final answers.
 As for TransferNet, the major weakness is the precision of explanations.
-As analyzed above, TransferNet tends to return redundant relations and, consequently, triples in final explanations.
-This is also the reason that it performed better than GNN2R on PQ-3hop regarding recall.
-However, a higher recall does not necessarily signify a better quality of explanation, especially given its lower precision.
+This is consistent with the above analysis that TransferNet tends to return too many relations and, consequently, redundant triples in final explanations.
+This contributed to a higher recall, especially on PQ-3hop.
+However, given its lower precision, the higher recall cannot signify a better quality of explanation.
 VI Conclusion